OpenStudy (iwanttogotostanford):

HELP PLZ REALLY SIMPLE BUT FORGOT: 17 divided by 4 and 1/4

OpenStudy (alexandervonhumboldt2):

\[\frac{ 17 }{ 4 }/\frac{ 1 }{ 4 }=\frac{ 17 }{ 4 }*4=\frac{ 17 }{ \cancel 4 }* \cancel 4=?\]

OpenStudy (md152727):

@AlexandervonHumboldt2

OpenStudy (iwanttogotostanford):

17?